Compare the monthly payments and total loan costs for the following pairs of loan options. Assume that both loans are fixed rate and have the same closing costs.

You need a ​$ 60,000 loan.

Option​ 1: a​ 30-year loan at an APR of 6.15%

Option​ 2: a​ 15-year loan at an APR of 5.75​%.

1-The monthly payment for option 1 is:

2- the monthly payment for option 2 is:

3- loan cost for option 1 is:

4- loan coast for option 2 is:

5- which loan option is the best:
Answer #1

Formula to Calculate EMI = (P x (r/12) x(1+(r/12))^n*12)/)((1+(r/12))^n*12)-1)

Ans 1) The monthly payment for option 1 is 365.54

=(60000*(0.0615/12)*(1+(0.0615/12))^360)/(((1+(0.0615/12))^360)-1) = 365.54

Ans 2) The monthly payment for option 2 is 498.24

=(60000*(0.0575/12)*(1+(0.0575/12))^180)/(((1+(0.0575/12))^180)-1) - 498.24

Ans 3) Loan cost for option 1 is 131593.3

=(P*(r/12)*n*12)/(1-(1+(r/12))^-n*12)

=(60000*(0.0615/12)*360)/(1-(1+(0.0615/12))^-360) = 131593.3

Ans 4) Loan cost for option 2 is 89684.29

=(60000*(0.0575/12)*180)/(1-(1+(0.0575/12))^-180) = 89684.29

Ans 5) Which loan option is the best:- Option 2

In option 2 loan cost is less in comparison to option 1 so option 2 is better though EMI outflow is little higher there but cost wise it's very effective.

Hope this helps. Feel free to share your feedback. Thanks and have a good day.